Let L be the line given by (3,-1,2) + t (1,1,-1).Show that the above line L lies on the plane -2x + 3y -4z + 1 = 0


1
Expert's answer
2021-07-13T12:50:25-0400
"<3,\u22121,2>+t<1,1,\u22121>"

"x=3+t,y=\u22121+t,\nz=2\u2212t"

Show that the Line lies on the plane 


"-2(3+t) + 3(-1+t) -4(2-t)+ 1 = 0"

"-6-2t-3+3t-8+4t+1=0"

"5t=16"


"t=3.2"

"x=6.2,y=2.2,z=-1.2"

Let "t=0"


"x=3,y=-1,z=2"

Check whether the point "(3, -1, 2)" lies on the plane


"-2(3)+3(-1)+4(2)+1=0, True"

Therefore the line L lies on the plane -2x + 3y -4z + 1 = 0.